When the standard deviation used to construct a confidence interval for a population mean increases the margin of error of the confidence interval to



A.Decreased



B.Constant



C.Increased



D.Remains the same

1
Expert's answer
2021-11-09T00:03:09-0500

The margin of error of the confidence interval for a population mean


"Margin\ of\ error=critical\ value\times\dfrac{SD}{\sqrt{n}}"

When the standard deviation used to construct a confidence interval for a population mean increases the margin of error of the confidence interval to

C.Increased